OpenStudy (anonymous):

Solve the system by using a table. -0.2x - 0.3y = 0.2 0.4 - 0.3y = 1.4

OpenStudy (phi):

Hero tried to help with this. It would help if you fix the 2nd equation.

OpenStudy (anonymous):

i still have no idea how to do it.

OpenStudy (anonymous):

here is the solution first of all get rid of all decimals (-2x-3y)/10=2/10 = -2x-3y=2 (4x-3y)/10=14/10 =4x-3y=14 4(-2x-3y)=2x2 2(4x-3y)=14x2 -8x-12y=4 8x-6y=28 8x cancels out so -18y=32 y=-32/18 y=-16/9 put y=-16/9 in -2x-3y=2 -2x+3*(16/9)=2 -2x+16/3=2 -6x+16=6 -6x=6-16 -6x=-10 x=10/6 x=5/3 answer is x=5/3 and y= -16/9

OpenStudy (anonymous):

these are the options i have (2, 2) (2, –2) (2, 1) (–2, 2)

OpenStudy (anonymous):

in second equation is it 0.4 or 0.4x

OpenStudy (anonymous):

0.4x